hi
I would like to know if I solved this equation correctly:
3x+y=25 where x-y=7 what is the value of y.
This is how I answered;
y=7-x
3x+ (7-x)=25
2x+7=25
2x+7=25
2x=18
x=9
y=7-9
y=-2
(x,y)= (9,-2)










Answer by Alan3354(69443)

hi
I would like to know if I solved this equation correctly:
3x+y=25 where x-y=7 what is the value of y.
This is how I answered;
y=7-x
3x+ (7-x)=25
2x+7=25
2x+7=25
2x=18
x=9
y=7-9
y=-2
(x,y)= (9,-2)
---------------
You don't need to ask, just sub your answer into the equations.
----------
3x+y=25
3*9 - 2 = 25 That works.
-------------
x-y=7
9 - (-2) = 7
11 = 7 uhoh. It's not correct.